calculate the static and dynamic lung pressure. peak pressure 40 cmH2O plateu preasure 32 cm H2O peep 5 cmH2O exhaled volume 400 ml
Answer #1

Lung pressure or pulmonary compliance is the ability of lung to stretch or distend.

Lung pressure or compliance is calculated by the general formula,

Compliance = Change in lung volume/ Change in pressure gradient during breathing

Calculation of static lung pressure:

Static lung pressure or static lung compliance is the pressure of the lung when there is no movement of air. It is given by the following formula,

Static lung pressure = VT/(Pplat - PEEP)

where, VT = Tidal volume or the amount of air displaced from lungs = 400 ml = 0.4 L

Pplat = Plateu pressure or the pressure applied by the alveoli = 32 cm H2O

PEEP = Positive end expiration pressure(pressure in the lung that is more than atmospheric pressure) = 5 cm H2O

Therefore, static lung pressure = 0.4 L/(32 - 5) cm H2O = 0.4/27 L/cm H2O = 0.015 L/cm H2O.

Calculation of Dyanmic lung pressure:

Dynamic lung pressure is defined as the change in volume of lungs during active movement of air in the lungs.

It is given by the following formula,

Dynamic lung pressure = VT/(PIP - PEEP)

where, VT = Tidal volume or the amount of air displaced from lungs = 400 ml = 0.4 L

PIP = Peak inspiratory pressure applied by the lungs = 40 cm H2O

PEEP = Positive end expiration pressure(pressure in the lung that is more than atmospheric pressure) = 5 cm H2O

Therefore, dynamic lung pressure = 0.4 L/(40 - 5) cm H2O = 0.4/35 L/cm H2O = 0.0114 L/cm H2O.